OpenStudy (anonymous):

how to solve an equation in quadratic form 3d^2/3-d^1/3-4=0 MEDAL AND FAN WILL BE GIVEN

OpenStudy (anonymous):

@dan815 hey can you help

OpenStudy (anonymous):

i know how to start it js

OpenStudy (jdoe0001):

the simpler way would be to simply multiply both left and right sides by 3 thus getting rid of the denominators

OpenStudy (anonymous):

do you know the method when you replace the common with u ?

OpenStudy (anonymous):

like 3(d^1/3)^2- d(1/3)-4=0 let (d^1/3) = u

OpenStudy (anonymous):

and then it would be 3u^2-u-4=0

OpenStudy (jdoe0001):

hmm

OpenStudy (anonymous):

is there a sum and product for -1 and -12

OpenStudy (jdoe0001):

3-4 = -1 3* -4 = -12

OpenStudy (anonymous):

thank you!!!! @jdoe0001

OpenStudy (jdoe0001):

yw